Which ordered pair would be the solution to this system of equations?

3x-4y=20
y=-3x+10

A(4,2) B(4,-2) C(-4,-2) D-4,22)
1
Expert's answer
2019-05-21T09:07:23-0400


Check if the first pair is a solution of the system:

A(4,2) 

"\begin{alignedat}{2}
 3*4- 4*2=4 {=}\mathllap{/\,} 20 \\
 2{=}\mathllap{/\,}-3*4 +10=-2
\end{alignedat}"


A(4,2) is not the solution to this system of equations

B(4,-2)

"\begin{alignedat}{2}
 3*4- 4*(-2)=20 \\
 -2=-3*4 +10
\end{alignedat}"

B(4,-2) is the solution to this system of equations

C(-4,-2)

"\begin{alignedat}{2}
 3*(-4)- 4*(-2)=-4 {=}\mathllap{/\,} 20 \\
 -2{=}\mathllap{/\,}-3*(-4) +10=22
\end{alignedat}"

C(-4,-2) is not the solution to this system of equations

D(-4,22)

"\begin{alignedat}{2}
 3*(-4)- 4*22=-100 {=}\mathllap{/\,} 20 \\
 22{=}\mathllap{/\,}-3*(-4) +10=22
\end{alignedat}"

D(-4,22) is not the solution to this system of equations

ANSWER:

B(4,-2) is the solution to this system of equations